Mars has an atmosphere composed almost entirely of carbon dioxide, with an average temperature of -63 ?C.

Part A

What is the rms speed of a molecule in Mars's atmosphere? Assume R = 8.315 J/mol?K.

Express your answer to two significant figures and include the appropriate units.
Answer #1

Speed \((\mathrm{rms})=\) sqrt \(3 \mathrm{KT} / \mathrm{m}\)

Where:

\(\mathrm{K}=\) Boltzmann's constant \(=1.38 \times 10^{-23}\) joule \(/ \mathrm{K}\)

\(\mathrm{T}=\) temperature in Kelvin degrees \(=273-63=210 \mathrm{~K}\)

\(\mathrm{m}=44^{*} 1.672 \times 10^{-27} \mathrm{Kg}\)

Molecular Speed \((\mathrm{CO} 2)=\operatorname{sqrt} 3^{*} 1.38 \times 10^{-23} * 210 /\left[44^{*} 1.672 \times 10^{-27}\right]\)

Speed \(=343.76 \mathrm{~m} / \mathrm{sec}\)